A body is moving in a circle with a uniform speed v,angular frequencw,time period T and frequency f.The centripetal acceleration can be given by:
(a)4pi2r/T
(b)v2/r
(c)wv
(d)wf
(e)4pi2rv2
(f)4pi2rf2
(PLEASE NOTE THAT THIS IS A MULTIPLE ANSWER QUESTION)
 

Grade:10

1 Answers

Aarti Gupta
9 years ago
Centripetal acceleration – Whenever an object is undergoing uniform circular motion, acceleration acting on it is called as centripetal acceleration.This acceleration always acts on the object along the radius towards the centre of the circular path.
This is given as -
ac = v2/ r which is option (b)
also,
v = w*r
putting we get,
ac = v * w which is option ( c)
Thus answers will be option b and option c.
ac = v*w